A pack of 25 kg with wooden roofs in 1 man has a 35 slope
pushing it to the roof. Between previously installed wooden roof coverings and the package
The kinetic friction coefficient is 0.32. 0.2 m / s2 of the package
up the roof with the acceleration
Find the force required to push it. (Take g = 9.8 m / s2.)
1
Expert's answer
2020-08-19T09:49:29-0400

"ma=F-mg\sin{35}-\mu mg \cos{35}\\25(0.2)=F-(25)(9.8)\sin{35}\\-0.32 (25)(9.8) \cos{35}"

"F=210\ N"